40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 190 of 348  Not logged in ELOG logo
ID Date Author Type Category Subject
  7962   Wed Jan 30 11:18:31 2013 ManasaUpdateScatteringScattering setup

Quote:

 [Jan, Manasa]

We are trying to get some scattering measurements in the Y-arm cavity. We have removed one of  the viewport windows window covers of ETMY chamber and have installed cameras on a ring that clamps to the window. The window along with the ring attachment is covered with aluminium foil when not in use.

[Jan, Manasa]

To align the camera to see small angle scattering from the ITMY, we tried shooting a green laser pointer at the pickoff mirror that was installed in the ETMY chamber such that we hit the face of ITMY. But we concluded that to be a very bad way to align the camera because we have no means to reconfirm that the camera was exactly looking at the scattering from ITMY.

Since we are in air, we came up with a plan B. The plan is to temporarily install a mirror in the ITMY chamber to steer the beam from the laser pointer (installed on the POY table) through ITMY to the pickoff mirror at the ETMY end. This way, we can install the camera at the ETMY window and be sure we are looking at ITMY scattered light. 

  7961   Wed Jan 30 11:16:32 2013 JenneUpdateLockingMode spacing calc

[Jenne, Jamie]

We did a few pen and paper calculations yesterday to confirm for ourselves that the half PRC should have nicely separated modes.  The half cavity is L=4.34m long, assuming flat mirror is 3.5 inches in front of BS.  That 3.5" is a guess, not a measurement.

Finesse

F = ( pi * sqrt(r1 * r2) ) / (1 - r1*r2) = 111.

Full width at half max

FWHM = c / (2 * L * F) = 311 kHz

FWHM in meters = FWHM * L/f = L*1064nm/c = 4.8 nm

Free spectral range

nu_fsr = F * FWHM = 34.5 MHz

Mode Spacing (eq 19.23 from Siegman)

omega = (n + m) * arccos(\pm sqrt(g1*g2)) / pi     *   (2*pi*c)/(2L)

For our half cavity, g1*g2 = 0.96

For the 01 or 10 modes, n+m = 1

omega = 13.7e6 rad/sec

mode spacing between 00 and 01 = 2.2 MHz

Thus, the modes should be well separated

=>  spacing is 2.2 MHz while FWHM is 0.311 MHz  (cavity fsr = 34.5 MHz)

 

EDIT JCD 31Jan2013:  Fixed mode spacing eqn to be diff between TEM00 mode and HOM, not plane wave and HOM.  Then fixed the factor of 2 error in the mode spacing numbers.

  7960   Wed Jan 30 03:01:55 2013 KojiUpdateGeneralEarly work on Mirror Mounts

I can't believe that SR785 can have such a low input noise level (<1nV/rtHz). Review your calibration again.

It is also described in the manual that SR560 typically has the input noise level of 4nV/rtHz, although this number depends on which gain you use.

  7959   Tue Jan 29 21:07:48 2013 JenneUpdateLockingPRM coils need diagonalizing

 

 [Jenne, Jamie]

We tried actuating on PRM so that we go through fringes in a known, linear way.  We used C1:SUS-PRM_LSC_EXC and awggui.  It seems that we get a lot of angular motion when we actuate....we need to look into this tomorrow.

EDIT/UPDATE:  Last night we tried several combinations of frequency and amplitude, but just for an idea,  we were using 2Hz, 1000cts.  Using Kiwamu's calibration in elog 5583 for the PRM actuator of 2e-8/f^2 m/cts, this means that we were pushing ~5nm.  But when we pushed much harder (larger amplitude) than that, we saw angular fringing. 

  7958   Tue Jan 29 20:28:11 2013 ericqUpdateGeneralEarly work on Mirror Mounts

 [Q, Chloe]

Chloe has been to the lab twice to start up her investigations in acoustic noise coupling to mirrors. The general idea for the setup is a HeNe laser bouncing off a mirror and onto a QPD, whose signal provides a measure of beam displacement noise. The mirror will be mounted and excited in various ways to make quantitative conclusions about the quality of different mounting schemes.

We have set up the laser+mirror+QPD on the SP table, and collected data via SR560s->SR785, with the main aim of evaluating the suitability of this setup. The data we collected is not calibrated to any meaningful units (yet). For now, we are just using QPD volts.

Chloe collected data of vertical displacement noise for the following schemes: Terminated SR785 input, Terminated SR560 inputs, Laser centered directly onto the QPD, Laser shining on mirror centered on QPD, laser/mirror/qpd with some small desktop speakers producing white noise from http://www.simplynoise.com. Data shown below. 

early.pdf

 

 

  7957   Tue Jan 29 19:50:49 2013 JenneUpdateLockingBetter POP layout, no extra PRM motion with locked cavity

[Jenne, Jamie, Manasa]

Today's activities focused on getting the POP layout improved, so that we could get clean data for the mode scan measurement. 

As Jamie and Koji pointed out yesterday, the beam was still a little too big on the POP DC PD, and was falling off the diode when the beam moved a small amount.  We have fixed things so that the PD is now at the focus of the lens, and the camera is at a place where the beam takes up most of the area on the TVs.  The beam no longer falls off the PD with cavity fluctuations.  A key point of this work was also to use an extra 2" optic to steer the beam down the length of the POP table, and then do the 50/50 beam splitting later with a 1" optic.  The 1" BS that we had been using (including with the "real" POP beam) is too small.  We could not find a 2" 50/50 BS, so we opted to do the splitting closer to the focal point.  Also, the BS that was splitting the beam between the PD and the camera was a 33% reflector, but now is a 50/50 BS. When we put back the 'real' POP path, we need to consider using larger optics, or a faster lens. The POP path is now good, hopefully for the duration of the half cavity test.

After getting the POP path taken care of, and tweaking up the cavity alignment a little bit, the transmitted power on POP DC is ~22,000 counts, with occasional fluctuations as high as 25,000 counts. 

Jamie looked at the REFL path, and things look sensible there.  The unlocked REFL power is ~36 counts, and the locked power is ~20 counts.  I'm not sure what the 160 counts that Koji mentioned in his edits to elog 7949 is about.

I looked at the PRM oplev with the cavity locked and unlocked, and with today's alignment, there seems to be no difference in the amount of PRM motion when the cavity is locked vs unlocked. 

HalfPRCL_PRM-flatMirror_RefsAreLocked_OthersUnlocked.png

 It still looks like we might be seeing some clipping in the in-vac POP steering mirrors - we haven't gotten to them yet.

Jamie is currently modifying Yuta's mode scan analysis script to look at the data that we have of the cavity.

 


We need more 2" optics.  There are no mounted 2" spares in the various optic "graveyards" (which, PS, we should consolidate all into the cabinet with doors near the optics bench), and the options for boxes in the drawers is slim pickin's.  We have some S-pol stuff, but no Y1s or BS-50s for P-pol.  Since POP, POX, POY, IPANG, TRX and TRY all come out of the vacuum with large beams, we should have some options for these laying around for this kind occasional temporary thing.  We also need to choose, then purchase better 2" lenses for the pickoffs.

Attachment 1: HalfPRCL_PRM-flatMirror_RefsAreLocked_OthersUnlocked.pdf
HalfPRCL_PRM-flatMirror_RefsAreLocked_OthersUnlocked.pdf
  7956   Tue Jan 29 18:40:20 2013 RijuUpdate Photodiode transimpedance

Today I have taken data for shot noise intercept current for PDA10CF. I will process the data and report.

Note: GPIB address changed, new command for AG4395A network/spectrum analyzer: ./netgpibdata.py -i 192.168.113.108 -d AG4395A -a 10 -f filename

  7955   Tue Jan 29 15:16:18 2013 ManasaFrogsGeneraltrial run

 

 I would like to suggest a trial run on these....Ergomates and the cleanboot!

ErgoMatesFeatures.jpg

booties.jpg

REPLY by JCD:  Are these going to trap dirt and be impossible to clean though?  The nice thing about Crocs and the giant flip flops is that they are solid and if they get dirty you can do a quick wipedown, and they're good as new.

  7954   Tue Jan 29 14:34:42 2013 JenneUpdatePEMSecret Seis filters

 

The BLRMS have been bad again, since the computer crash of last week.  Finally getting around to looking into it, I discovered that there are filter banks that have the microns/sec calibration filters, which are not accessible from the sitemap.  I have added links to them for GUR1 and GUR2.  We need to make the PEM/BLRMS screens macro-expansion-y, so that I don't have to change each screen individually.

Anyhow, the BLRMS are back.

 

  7953   Tue Jan 29 14:20:02 2013 KojiUpdateGeneralFiner rotation stage for optics characterization

A rotation stage has been ordered.

Newport Rotation Stage, 360° Coarse, 5° Fine Rotation, Micrometer  Newport 481-A
Newport Solid Insert for RSP-1T Rotation Stage Newport RSA-1TI
Newport Universal Mounting Plate, 2.56 in. x 2.56 in. x 0.5 in., 1/4-20 Thread  Newport UP-1A

Specification: Newport 481-A

  • Sensitivity: 15 arcsec
  • Graduations: 1 deg
  • Vernier: 5 arcmin
  • Fine travel range: 5 deg
  • With Micrometer
  7952   Tue Jan 29 10:59:37 2013 lazy personFrogsGeneralbetter plan

 

 I propose we work around this problem with giant flip-flops.  These are in the vein of the take-off-your-shoes-and-put-on-Crocs, without the taking off your shoes part.  They're a little annoying on the sticky mats, but otherwise great.  They are also super easy to put on and take off without hands, so there's no excuse for wearing them around the control room. 

I propose we buy many pairs of the smalls in green (since we already have one green small...they are big on me, so should be just right for most people), and a few mediums in, say, blue, and a few larges in black, and then maybe a few extra larges in green for people with extraordinarily large feet (they only have 3 colors).  Then we can keep a few pairs of each by each door to the lab, and have no more tracking dirty control room filth into the lab.

  7951   Tue Jan 29 10:50:02 2013 JenneUpdateLockingVideo of PRM-flat test cavity

 

I think Den accidentally edited and overwrote my entry, rather than replying, so I'm going to recreate it from memory:

I aligned the PRM-flat test cavity (although not as well as Jamie and Koji did later in the evening) and took some videos. Note that these may not be as relevant any more, since Jamie and Koji improved things after I left.

 

Also, before doing anything with the cavity, I tuned up the PMC since the pitch input alignment wasn't perfect (we were getting ~0.7 transmission), and also tuned up the MC alignment and remeasured the MC spot positions, to maintain a record.

  7950   Mon Jan 28 21:36:44 2013 tall guyFrogsGeneralsmall people on notice

If I catch anyone putting small booties into the large bootie bin, I will make said person eat small booties.

  7949   Mon Jan 28 21:32:38 2013 jamieUpdateAlignmenttweaking of alignment into half PRC

[Koji, Jamie]

We tweaked up the alignment of the half PRC a bit.  Koji started by looking at the REFL and POP DC powers as a function of TT2 and PRM alignment. 
He found that the reflected beam for good PRC transmission was not well overlapped at REFL.  When the beam was well overlapped at REFL, there was clipping in the REFL path on the AS table.

We started by getting good overlap at REFL, and then went to the AS table to tweak up all the beams on the REFL pds and cameras.
This made the unlocked REFL DC about 40 count. This was about 10mV (=0.2mA) at the REFL55 PD.
This amazed Koji since we found the REFL DC (of the day) of 160 as the maximum of the day for a particular combination of the PRM Pitch and TT2 Pitch. So something wrong could be somewhere.

We then moved to the ITMX table where we cleaned up the POP path.  We noticed that the lens in the POP path is a little slow, so the beam is too big on the POP PD and on the POP camera (and on the camera pick-off mirror as well)
We moved the currently unused POP55 and POP22/110 RFPDs out of the way so we could move the POP RF PD and camera back closer to the focus.  Things are better, but we still need to get a better focus, particularly on the POP PD.

We found two irides on the oplev path. They are too big and one of these is too close to the POP beam. Since it does not make sense too to have two irides in vicinity, we pulled out that one from the post.

Other things we noticed:

  • The POP beam is definitely clipping in the vacuum, looks like on two sides.
  • We can probably get better layout on the POP table, so we're not hitting mirrors at oblique angles and can get beams on grid paths.

After the alignment work on the tables, we started locking the cavity. We already saw the improvement of the POPDC power from 1000 cnt to 2500 cnt without any realignment.
Once PRM is tweaked a little (0.01ish for pitch and yaw), the maximum POPDC of 6000 was achieved. But still the POP camera shows non-gaussian shape of the beam and the Faraday camera shows bright
scattering of the beam. It seems that the scattering at the Faraday is not from the main beam but the halo leaking from the cavity (i.e. unlocking of the cavity made the scattering disappeared)


Tomorrow Jenne and I will go into BS to tweak the alignment of the TEMP PRC flat mirror, and into ITMX to see if we can clean up the POP path.

  7948   Mon Jan 28 19:15:14 2013 ManasaUpdateScatteringScattering setup

 [Jan, Manasa]

We are trying to get some scattering measurements in the Y-arm cavity. We have removed one of  the viewport windows window covers of ETMY chamber and have installed cameras on a ring that clamps to the window. The window along with the ring attachment is covered with aluminium foil when not in use.

  7947   Mon Jan 28 19:07:45 2013 ManasaUpdateGeneralSN6 Laseroptik mirror - Tranmittance measurements

I repeated the measurements using NPRO instead of Crystalaser. I am attaching optical layouts for these measurements for future reference. 

Lesson learnt : Do not use Crystalaser for transmission measurements and always separate the transmitted main beam from other beams that result from the wedged surface of the mirror. 

trans_layout.png

Measurements match specs provided by Laseroptik

p-polarization

T percentage = 0.10% 42 deg
                            0.092% 44 deg
                            0.086% 46 deg
Minimum transmittance = 0.081% 52deg

s-polarization

T percentage = 0.048% 42 deg
                            0.047% 44 deg
                            0.047% 46 deg
Minimum transmittance = 0.047% 46 deg

sn6_trans0128.png
 

  7946   Mon Jan 28 17:59:02 2013 RijuUpdate Photodiode transimpedance

Summary: Measurement and plot of shot-noise-intercept-current for PDA255.

Motivation:It is to measure the shot noise intercept current for PDA255 - the MC transmission RF photodiode to get an idea for the noise current for the detector

Result: The final plot is attached here. The plot suggests that the value of shot-noise-intercept current is 3.06mA

Discussion:

The plot is for the measured data of Noise voltage (V/sqrt(Hz)) vs DCcurrent(A). The fitted plot to this measured data follows the noise equation

Vnoise = gdet* sqrt[ 2e (iDC+idet)] ,  where gdet= transimpedance of the PD in RF region as described in manual of PDA255 (i.e. 5e3 when it is not in High-impedance region).

On the other hand for DCcurrent calculation we must use the high-impedance value for the transimpedance i.e. 1e4 Ohm. idet is the shot noise intercept current.

For the rough calculation of the noise level we may use the following formulae:

Vnoise = gdet*sqrt[2e (iDC+idet)] = gdet*sqrt(2e in), when in=iDC+idet;

For say, in1=1mA; Vnoise1=gdet*sqrt(2e *in1)

and sqrt(2e *in1)~18pA/sqrt(Hz)

In current case dark noise is ~1.5e-7 V/sqrt(Hz)

Therefore dark current(in2) ~dark noise voltage/RF transimpedance = 30pA/sqrt(Hz)

i.e. sqrt(2e *in2)=30pA/sqrt(Hz)

i.e. sqrt(in2/in1)=30/18

therefore, in2~3mA (since in1=1mA)

For, iDC=0, in=idet.

Therefore the shot-noise-intercept current will be ~3mA

Then Vdc = in2*1e4 = 30V

According to the experiment  and also from the PDA255 manual the DC voltage level never goes beyond ~10V. Therefore following the photodiode characteristics(we work in reverse bias) we may infer that it can never become shot noise limited.

Also, from PDA255 manual, at 1650nm the dark noise is 30pW/sqrt(Hz) and the responsivity is 0.9A/W. Therefore the noise current level will be = noise power* responsivity ~27pA/sqrt(Hz). The value matches well with our expectation.

 

Attachment 1: shotnoiseinterceptpda255.pdf
shotnoiseinterceptpda255.pdf
  7945   Mon Jan 28 17:01:19 2013 DenUpdateLockingVideo of PRM-flat test cavity

What mode will you get if lock the cavity PRM - ITMY/ITMX/TEST MIRROR without PR2, PR3 and BS?

Is it possible to skip MC1, MC3 and lock the laser to this test cavity to make sure that this is not actuator/electronics noise?

  7944   Fri Jan 25 08:10:42 2013 SteveUpdateGeneralLaseroptik mirrors

 Here are the German plots. Unfortunately they are not logarithmic.

 Ed: Proprietary data removed. Use wiki (Koji)

Steve uploaded data to the 40m wiki  / Aux_Optics on 02-07-2013

 

 

  7943   Thu Jan 24 16:34:56 2013 ranaUpdateGeneralLaseroptik mirror - SN6

  I have two questions:

1) Are we sure that the T measurement is not being compromised by some systematic? i.e. some leakage is making the apparent T appear too high.

2) IF the T is really so high, how should we decide whether or not to use this one rather than the G&H? Is the 532 nm property more important than the high recycling gain?

  7942   Thu Jan 24 16:31:46 2013 SteveUpdatePEMbuilding exterier wall painted

The wood exteier walls, gutters and doors were painted at CES-Annex building #69

Attachment 1: IMG_1880.JPG
IMG_1880.JPG
  7941   Thu Jan 24 16:23:24 2013 KojiUpdateGeneralLaseroptik mirror - SN6

The mirror T is completely out of spec. We should find or request the data sheet of the mirror.

> We should consider buying a better rotation stage

I'm already on it

  7940   Thu Jan 24 15:16:50 2013 ManasaUpdateGeneralLaseroptik mirror - SN6

I repeated the transmittance measurements of Laseroptik SN6 at 1064nm. The rotation stage could only resolve 2 deg rotation (We should consider buying a better rotation stage).

s-polarization

Percentage transmittance

       0.177% 42 deg
       0.806% 44 deg
       0.57%   46 deg
       54.8%    0 deg

p-polarization

Percentage transmittance

       1.039% 42deg
       1.155% 44 deg
       1.159% 46 deg
        65.6%   0 deg

sn6_trans1.png 

  7939   Thu Jan 24 14:40:09 2013 SteveUpdateVACRGA backgroung

 RGA background with VM2 open to Maglev at day 37

 

Note: The PAN gauge of the  annulos is at atm.  Please do not vent this 200 ft long annulos line when you venting the annulos of a chamber. The chamber annulos should be closed off to this long 2"  OD. pipe before you vent the annulos of a chamber.

 

Attachment 1: Screenshot.png
Screenshot.png
  7938   Thu Jan 24 11:40:55 2013 JenneUpdateSUSETMX damping restored

Quote:

ETMX sus damping restored.  It is still noisy

 I should have elogged, but I turned off the watchdog to remind myself that iscex computer is still crashed.  "Turning on" the damping doesn't do anything since there aren't any signals going to the coils from the computer.

  7937   Thu Jan 24 11:31:45 2013 SteveUpdateSUSETMX damping restored

ETMX sus damping restored.  It is still noisy

  7936   Thu Jan 24 08:28:35 2013 SteveUpdateVACwhen to pumpdown ?

Reminder: I will be on vacation next week. We would have to put the access connector in tonight if I'm pumping down Friday, tomorrow.

  7935   Wed Jan 23 22:02:25 2013 ManasaUpdateGeneralLaseroptik mirror - SN6

Quote:

Got confused (even after I talked with Manasa).

The plot shows the number ~0.01 or less at 45deg. But the number is the text does not match with the plot.

Please use the logarithmic scale for the vertical axis.
And more points between 35 to 50 deg please (like ~1deg spacing)

Don't we have the data sheet from the coater? Can we request it?

I corrected the typo in the text...however, I agree the plot was lame...Will get the data sheet made tomorrow! 

  7934   Wed Jan 23 20:46:46 2013 Zen MasterUpdateLockingPR-flat cavity status - locks!

Quote:

I (with help from Q)

 Two quadratures working in harmony.

yinYang.png

  7933   Wed Jan 23 20:27:05 2013 RijuUpdate Photodiode transimpedance

Today I have repeated the expt for shot noise intercept current. Koji found that the Spectrum analyzer is going to saturation, so we have used one DC blocker (MCL - 15542 model) in PD signal.

I will analyze the data and report.

Ed by Koji: DC BLOCK is  BLK-89-S

  7932   Wed Jan 23 20:24:05 2013 KojiUpdateGeneralLaseroptik mirror - SN6

Got confused (even after I talked with Manasa).

The plot shows the number ~0.01 or less at 45deg. But the number is the text does not match with the plot.

Please use the logarithmic scale for the vertical axis.
And more points between 35 to 50 deg please (like ~1deg spacing)

Don't we have the data sheet from the coater? Can we request it?

  7931   Wed Jan 23 19:05:16 2013 JenneUpdateLockingPR-flat cavity status - locks!

Status update

I (with help from Q) have redone the POP path on the ITMX table.  1" iris is a little too small, so I took it out.  2" lens moved to be centered on POP beam.  2" Y1 didn't need moving.  Straight refl from the 2" Y1 was aligned on to a PDA10CS (set to 70dB). This PD is blocking the usual POP55 diode.  BS which sends beam to camera was moved to allow room for the new temp DC PD.  Refl from this BS goes to the POP camera, which was moved so that the POP beam takes up most of the camera.  BS that would normally take half of the camera's beam and send it to POP22 (Thorlabs PD) is removed, so no beam to POP22.

Also, I have taken the output of the PDA10CS and hijacked the "POP110" heliax cable.  This was connected to this Thorlabs PD which is used as POP22.  (Kiwamu and I had long-term borrowed the 110 demod board for an AS 110 diode, so the "POP110" heliax was really only serving POP22.) There are yellow labels on the new temp and old regular cables, so we can undo my hack.  Similarly, on the other end of the heliax at the LSC rack, I have taken the heliax's output and sent it to the POPDC input on the whitening board.  Thus, the regular POPDC SMA cable is unplugged, but labeled again with big yellow labels.

In other news - the PR-flat cavity locks!!! 

Koji and I coarsely rotated the REFL11 phase such that the signal is predominantly in the I phase.  We set the LSC input matrix to use REFL11I for PRCL, and the output matrix is set to actuate on PRM.  Then we set the gain to -0.005, and it locked!!!!

 

EDIT:  I turned back on the PRM oplev (after Manasa aligned it and redid the out-of-vac oplev layout a bit), and the motion of the cavity is slightly reduced, although there's still a lot going on.  The cavity is vaguely well aligned, although it's time to go make sure that the beams are still on the REFL and TRANS PDs.  However, it's dinner time.

  7930   Wed Jan 23 18:16:11 2013 ManasaUpdateGeneralLaseroptik mirror - SN6

 I repeated the transmittance measurements of LaserOptik SN6 @1064nm.

Transmittance for s-polarization 

0 deg - 0.524
45 deg - 0.055

Transmittance for p-polarization

0 deg - 0.515
45 deg - 0.1047 0.01047

Raji's measurements are here.

Attachment 1: sn6_trans.png
sn6_trans.png
  7929   Wed Jan 23 11:43:19 2013 RijuUpdate Photodiode transimpedance

Quote:

- The data should be plotted in a log-log scale.
- The data points were only taken in the high current region.

- The plot may suggest that the amplifier saturate at the RF.

PDA255 has the nomial transimpedance gain of 10^4 Ohm.
The DC current of 10^-3 gives the output of 10V.
This plot may tell that the saturation starts even at the 1/10 of the full DC range.

The plot doesn't have many points below 0.1mA.
Consult with my plots for the similar measurements.
The measured points are logarithmically spaced. Use the same technique.

- It is also very unknown that how the noise level is calculated. No info is supplied in the plot or the elogs.

 Here I am attaching the plot in loglog scale. I have taken the data-points from no light condition to the maximum light condition, the minimum variation possible in the current supply was 0.01A. The noise was visibly decreasing at higher light level.

For the noise level calculation I took the average of total noise in the range 7-60MHz. For each range the formula used was

noisevalue= sqrt(data(:,2)*100)/sqrt(2)/sqrt(channel BW);     -- this conversion is needed since the data was collected in the 2 column format: frequency, spectrum(W).

Attachment 1: pda255shotnoiseintercept1.pdf
pda255shotnoiseintercept1.pdf
  7928   Tue Jan 22 19:53:01 2013 JenneUpdateLockingPR-flat cavity status - not locked

The PR-flat cavity is flashing, although not locked.  I am too hungry to continue right now.

I put the FI_Back camera on a tripod, looking at the back of the Faraday.  The beam that Jamie and I were working with on Friday was clipped going back through the Faraday.  I twiddled the TT2 and PRM pointing such that the beam is retroreflecting, and getting back through the Faraday, and the cavity is still flashing.  I then redid the REFL path on the AS table a little bit.  The beam is currently going to the REFL camera, as well as REFL11 and REFL55. 

Some notes about the AS table:  The Y1 separating the main REFL beam from the REFL camera beam was mounted 90 degrees (rotated about the beam's axis) from what it should be.  I fixed it, so that the straight-through beam that goes to the camera is not clipped by the edge of the mount.  The reason (I think) this mirror was mounted backwards is that when mounted correctly, the back of the mount and the knobs interfere with the AS beam path.  I solved this by rotating the first out-of-vac REFL mirror a small amount so that the REFL and AS beams are slightly more separated. 

I am not seeing any nice PDH signal on dataviewer, so I went to check the signal path for the PDs.  The 11MHz marconi is on and providing RF, the EOM is plugged in to 11, 55 and 29.5 signals (no aux cavity scan cables are plugged in).  Both of the RF Alberto boxes are on.  I measured the RF output of both REFL11 and REFL55, although after the fact I realized that I was BAD, and had not found a 'scope that lets me change the input impedance to 50 ohms.  BAD grad student.  However, since I have numbers, I will post them, despite their being not quite correct:

284mVpp at 11MHz out of REFL11.  This is -6.9dBm

2mVpp at 55MHz out of REFL55, measured by 'scope

So, I can clearly see the 11MHz on the 'scope, and can see a very noisy, small 55MHz signal on the 'scope.  I need to think over dinner about what level of signal we should be sending to the demod boards, and whether or not I need more power coming out of the RFPDs.  There is a wave plate and PBS before beam goes to any of the REFL PDs, presumably to ensure that none of them get fried when we're at high power.  If I need more signal, I suspect I can rotate the wave plate and let more light go to the diodes.

  7927   Tue Jan 22 19:51:52 2013 KojiUpdate Photodiode transimpedance

- The data should be plotted in a log-log scale.
- The data points were only taken in the high current region.

- The plot may suggest that the amplifier saturate at the RF.

PDA255 has the nomial transimpedance gain of 10^4 Ohm.
The DC current of 10^-3 gives the output of 10V.
This plot may tell that the saturation starts even at the 1/10 of the full DC range.

The plot doesn't have many points below 0.1mA.
Consult with my plots for the similar measurements.
The measured points are logarithmically spaced. Use the same technique.

- It is also very unknown that how the noise level is calculated. No info is supplied in the plot or the elogs.

  7926   Tue Jan 22 17:29:29 2013 RijuUpdate Photodiode transimpedance

Riju

Summary:  I am stuck with the measurement of shot-noise-intercept-current of PDA255. Seeking help.

Motivation: It is to measure the shot noise intercept current for PDA255 - the MC transmission RF photodiode to get an idea for the noise current for the detector.

Method: It is as described in the elog  7907 

Result: The plot is attached here.

Discussion: The result I got is really unexpected, the noise voltage should increase with the DC current level that corresponds to the increment of light level too. But actually it is decreasing. Three times I have repeated this experiment and got the same result. I want some suggestion on this regard.

Attachment 1: pda255shotnoiseintercept.pdf
pda255shotnoiseintercept.pdf
  7925   Tue Jan 22 11:09:58 2013 SteveUpdateTip - TiltSS mirror holder ordered

Koji's design of the SS  2" mirror holder with flexure spring optic retainer  like Polaris-K1 has been ordered. We are getting just one to see it's effect on the hysteresis.

Attachment 1: D1201574_v1_Tip-Tilt_Mirror_Holder.SLDDRW
  7924   Tue Jan 22 09:18:50 2013 SteveUpdateTip - Tilttip tilt bases ordered

The corrected drawing base for tip tilt with coils are going to the shop. The will be back by the end of the week.

Attachment 1: TTbc.PDF
TTbc.PDF
  7923   Tue Jan 22 09:10:19 2013 SteveUpdatePEMbig foot is dirty

Please wipe, clean car wheels and wear booties entering the 40m lab.

Obviously this person has no idea about our clean room rules.

Attachment 1: IMG_1873.JPG
IMG_1873.JPG
Attachment 2: IMG_1874.JPG
IMG_1874.JPG
  7922   Sat Jan 19 18:23:31 2013 ranaUpdateComputersAll front ends but c1lsc are down

After sshing into several machines and doing 'sudo shutdown -r now', some of them came back and ran their processes.

After hitting the reset button on the RFM switch, their diagnostic lights came back. After restarting the Dolphin task on fb:

"sudo /etc/init.d/dis_networkmgr restart"

the Dolphin diagnostic lights came up green on the FE status screen.

iscex still wouldn't come up. The awgtpman tasks on there keep trying to start but then stop due to not finding ADCs.

 

Then power cycled the IO Chassis for EX and then awtpman log files changed, but still no green lights. Then tried a soft reboot on fb and now its not booting correctly.

Hardware lights are on, but I can't telnet into it. Tried power cycling it once or twice, but no luck.

Probably Jamie will have to hook up a keyboard and monitor to it, to find out why its not booting.

FE.png

P.S. The snapshot scripts in the yellow button don't work and the MEDM screen itself is missing the time/date string on the top.

  7921   Sat Jan 19 16:02:28 2013 ranaOmnistructureElectronicsPS cleanup

During our 'Women in Physics' tours today, we were reminded that there are several bench power supplies being used as permanent inside.

Some are being used to power PZTs, AOMs, VCOs, RFPDs, etc. On Wednesday, after the meeting, we will all go inside and remove one and replace it with a fused, professional wiring to the rack power supplies. The temporary ones must be removed.

  7920   Sat Jan 19 15:05:37 2013 JenneUpdateComputersAll front ends but c1lsc are down

Message I get from dmesg of c1sus's IOP:

[   44.372986] c1x02: Triggered the ADC
[   68.200063] c1x02: Channel Hopping Detected on one or more ADC modules !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!
[   68.200064] c1x02: Check GDSTP screen ADC status bits to id affected ADC modules
[   68.200065] c1x02: Code is exiting ..............
[   68.200066] c1x02: exiting from fe_code()

Right now, c1x02's max cpu indicator reads 73,000 micro seconds.  c1x05 is 4,300usec, and c1x01 seems totally fine, except that it has the 02xbad.

c1x02 has 0xbad (not 0x2bad).  All other models on c1sus, c1ioo, c1iscex and c1iscey all have 0x2bad.

Also, no models on those computers have 'heartbeats'.

C1x02 has "NO SYNC", but all other IOPs are fine.

I've tried rebooting c1sus, restarting the daqd process on fb, all to no avail.  I can ssh / ping all of the computers, but not get the models running.  Restarting the models also doesn't help.

upside_down_cat-t2.jpg

c1iscex's IOP dmesg:

[   38.626001] c1x01: Triggered the ADC
[   39.626001] c1x01: timeout 0 1000000
[   39.626001] c1x01: exiting from fe_code()

c1ioo's IOP has the same ADC channel hopping error as c1sus'.

 

  7919   Fri Jan 18 15:08:13 2013 jamieUpdateAlignmentalignment of temporary half PRC

[jenne, jamie]

Jenne and I got the half PRC flashing.  We could see flashes in the PRM and PR2 face cameras.

We took out the mirror in the REFL path on the AP that diverts the beam to the REFL RF pds so that we could get more light on the REFL camera.  Added an ND filter to the REFL camera so as not to saturate.

  7918   Fri Jan 18 12:08:08 2013 KojiUpdateLockingDust?

No

Quote:

My question is, could a few scattering centers cause the behavior that we are seeing?

 

  7917   Fri Jan 18 09:54:18 2013 JenneUpdateLockingPRM - Flat mirror cavity

Quote:

I do, however, see flashes on the face of the PRM.  Movie saved, will post soonly.

 Dang it.  I didn't confirm that the movie was good, just that it was there.  It's corrupted or something, and won't play.  I'll just have to make a new movie today after I realign the cavity.

  7916   Fri Jan 18 00:41:34 2013 JenneUpdateLockingDust?

I was thinking tonight about more possible reasons that our PRC sucks, and I wonder if dust on the BS could create the problem.

Historically, Kiwamu and I found a few dust particle scattering centers every time we inspected the test masses before drag wiping. Sometimes, there would be one frustratingly close to the center of the optic. I'm not sure if we ever made note of how many we saw and where they were, except out loud to the assembled crowd.

Anyhow, the BS is the only IFO optic that was not replaced, so I'm not sure how long it has been since it was cleaned. If the PR-flat cavity looks okay and we take out the BS to do a PRM-ITMY cavity, we should inspect the beam splitter.

Also, the PRM could need cleaning, but at least it has been drag wiped within recent memory.

My question is, could a few scattering centers cause the behavior that we are seeing?

 

EDIT:  List o' elogs....

 

Elog 5301 - Some details on dust seen on ITMs and ETMs, Aug 2011.

Elog 4084 - Kiwamu's in-situ drag wiping how-to, with details on some of the dust we saw. Dec 2010.

Elog 3736 - PRM drag wiped before suspension (Oct 2010)

Elog 3111 - June 2010, BS drag wiped.

  7915   Thu Jan 17 19:33:53 2013 ManasaUpdateAlignmentPRM oplev

 We had to work on redesigning the oplev layout in BSC when I found that the positions of the mirrors were clipping IPPOS and the green beam while updating the CAD layout.  

To avoid any clipping, the prm oplev beam is steered into the vacuum by an oplev mirror and out of vacuum through 3 steering mirrors. The table weights had to be moved to allow room for the oplev mirrors. Hence table had to be re-leveled. I will update the CAD drawing with the current position of the mirrors and will reconfirm that the new mirrors are not in the way of any of the beams. In-vac photos are updated in picasa.

  7914   Thu Jan 17 17:04:49 2013 SteveUpdateVACvac monitor screen changed

The COVAC_MONITOR.adl was changed. Ion pump labeled as disconnected means: ion pump controller power turned off and ion pump gate valve control connection disconnected.

RP2 roughing pump labeled disconnected means:  hardware  disconnected.

The actual operational, valve control screen has not been changed yet.

Attachment 1: screenmod.png
screenmod.png
  7913   Thu Jan 17 15:48:21 2013 JenneUpdateLockingPRM - Flat mirror cavity

 

 2" G&H mirror is installed on a DLC mount just in front of the BS.  I had to remove one of the 4 BS dog clamps, so we must put it back when we are finished with this test.

I aligned the G&H mirror such that the reflected beam is overlapped with the incident beam, and I aligned the PRM such that the regular REFL beam is retro-reflected.  This is the same as getting the beam bouncing off the PRM back to the G&H to be overlapped.

I then saw flashes of the cavity, when I held a card with a hole in the cavity, so the beam was going through a small aperture in the card, but I still saw flashes.  I was not able to see flashes on the IR card transmitted through the G&H mirror.

I also cannot see any flashes or scattered light on the face of PR2 camera.

I do, however, see flashes on the face of the PRM.  Movie saved, will post soonly.

Light is coming out of REFL on the AS table, but it's clipped somewhere....needs investigation/work before we can lock.

I also didn't see anything at the POP port with a card, but I'm hopeful that perhaps with a camera I'll see something.

ELOG V3.1.3-